OpenStudy (anonymous):

Could somebody give me a hand with the following question: Find an equation of the tangent line to the curve at the given point: y=sqrt(x), (49,7)

OpenStudy (anonymous):

\[ y'= \frac 1{2\sqrt x} \]

OpenStudy (anonymous):

Find the slope m=y'(49) and us y-y1 =m(x-x1)

OpenStudy (anonymous):

\[ y'(49)=\frac 1 { 2 \sqrt{49}}=\frac 1 {14} \]

OpenStudy (anonymous):

\[ y-7 =\frac 1 {14} (x-49) \]

OpenStudy (anonymous):

Thank you very much, I greatly appreciate your help, but I have another question to ask you if you don't mind.

OpenStudy (anonymous):

Since I'm just starting calculus, I have to take the long way around to get the solution; I have to use the equation for slope. I don't suppose you could show me how that works could you?

OpenStudy (anonymous):

Have they explained derivative to you?